OpenStudy (matt6288):

3^2+(-3)^2 explain or just tell me how to do (-3)^2 cuz its not -9?..

OpenStudy (pfenn1):

Remember\[(1)^2 = 1\]\[(-1)^2=1\]\[-(1)^2=-1\]

OpenStudy (matt6288):

ty :)

OpenStudy (dims116):

when you multiply negative with negative it gives u a positive sign. or any negative number to the power of any even number will give u a positive answer.

OpenStudy (pfenn1):

you're welcome